OpenStudy (tainted):

Complete each ordered pair so that it is a solution of the given linear equation. x-3y=2; (?,1) (-13,?)

OpenStudy (08surya):

5,-5

OpenStudy (gorv):

these point should satisfy the equation (?,1) ? this value ox we have to find and 1 is value of y so put it in equation and find the value of x

OpenStudy (gorv):

x-3*1=2 x-3=2 x=2+3 x=5

OpenStudy (gorv):

similarly calculate for other